You are deciding between two mutually exclusive investment opportunities. Both require the same initial investment of S10.4 million. Investment A will generate S2.09 million per year (starting at the end of the first year) in perpetuity. Investment B will generate $1.54 million at the end of the first year, and ts revenues will grow a 2.8% per year for every year after at. a. Which investment has the higher IRR? b. Which investment has the higher NPV when the cost of capital is 5.2%? c. In this case, when does picking the higher IRR give the correct answer as to which investment is the best opportunity? a. Which investment has the higher IRR? The IRR of investment A is | %. (Round to two decimal places.)
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard Alignment Number Cells Edting CO ca CR CS CT CU CV CX CY CZ DA DB DC INITIAL INVESTMENT INVESTMENT A INVESTMENT B 10.4 2.09 PERPETUAL 1.54 ANNUAL GROWTH OF 2.8% 67 69 70 71 72 73 74 75 76 IRR IS THE POINT WHERE NPV NPV PV OF ALL CASH INFLOWS PV OF ALL CASH OUTFLOWS NPV 2.09/(IRR) -10.4 0 2.09/(IRR) -10.4 10.4 2.09/(IRR) IRR = 20.10% IRR FORA NPV 1.54/(IRR - g)-10.4 0 1.54/(IRR-0.028) -10.4 10.4 = 1.54/( IRR-0.028) IRR-0.028 = 0.14808 IRR-17.61% IRR FOR B 78 79 80 81 82 83 84 11 BASED ON IRR, YOU WILL PICK INVESTMENT A AS IT HAS HIGHER IRR IRRnpv RATION PBP ARR ( external rate / MIRR NPV İRR REPLACEMENT S HPR GMAM EAC MACRS RATIOCASHBUDGET acc-BOND EPS EBIT- REPLACEMENT- STats- ト 福 130%Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy ▼ в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. ▼ ㆆ ▼ Clipboard Alignment Number Cells CR CS NPV FORA CO CT CU NPV (2.09/0.052) -10.4 NPV 40.1923 10.4 NPV = 29.7923 CV CX CY DA DB DC DE 86 87 89 90 91 92 93 94 95 96 97 98 NPV 1.54/(IRR - g) -10.4 NPV 1.54/(0.052-0.028) -10.4 NPV 53.7667 NPV FOR B NPV RULE REQUIRES YOU TO CHOOSE INVESTMENT B AS IT HAS HIGHER NPV HERE WE HAVE TO COMPARE BOTH INVESTMENT AND EQUATE THEM AND FIND OUT THE RATE 2.09/(RATE) -10.4 1.54/(RATE-g) 10.4 2.09/(RATE) 1.54/(RATE-0.028) 2.09 (RATE 0.028) 1.54 RATE 2.09 RATE-0.05852- 1.54 RATE RATE 10.64% 100 101 102 103 104 11 1 r THE BEST INVESTMENT OPPORTUNITY WHEN PICKING THE HIGHER IRR OCCURS FOR ALL DISCOUNTS RATES HIGHER THAN 10.64% IRRnpv RATION PBP ARR external rate MIRR NPV İRR REPLACEMENT S -HPR GMAM EAC MACRS RATIOCASHBUDGET wacc BOND | EPS EBIT- REPLACEMENT- STats | | |